OpenStudy (anonymous):

what is the sum of the following series: infinity E n=1 ((2^n+8^n)/11^n))

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \frac{ 2^n+8^n }{11^n}}\) like this?

OpenStudy (anonymous):

yes Im really lost ... :(

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \frac{ 2^n+8^n }{11^n}}\) \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \frac{ 2^n }{11^n}+\frac{ 8^n }{11^n}}\) \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \frac{ 2^n }{11^n}+\sum_{ n=1 }^{ \infty } \frac{ 8^n }{11^n}}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \left(\frac{ 2 }{11}\right)^n+\sum_{ n=1 }^{ \infty } ~ \left(\frac{ 8 }{11}\right)^n}\)

OpenStudy (solomonzelman):

Do you know about "geometric series" ?

OpenStudy (anonymous):

I think i understand it a bit. its =A/1-x?

OpenStudy (solomonzelman):

\(a_1~/~(1-r)\) where r is the common ratio and \(a_1\) is the first term of the series.

OpenStudy (solomonzelman):

Do you want an example?

OpenStudy (anonymous):

is the answer 1 then ?

OpenStudy (anonymous):

would it be: a= 10/11 the bottom would be 1-(10/11)

OpenStudy (solomonzelman):

lets do the first series, ok?

OpenStudy (anonymous):

yes, thank you.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \left(\frac{2 }{11}\right)^n}\) \(\large\color{black}{ \displaystyle a_1=2/11}\) So the sum is: \(\large\color{black}{ \displaystyle S=\frac{a_1 }{1-r}=\frac{\frac{2}{11}}{1-\frac{2}{11}}}\) can you simplify the sum further?

OpenStudy (anonymous):

so it comes to be (2/11)/(9/11)

OpenStudy (anonymous):

or 2/9

OpenStudy (solomonzelman):

yes, and that would simplify to?

OpenStudy (solomonzelman):

yes, 2/9 good

OpenStudy (solomonzelman):

Ok, can you do \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \left(\frac{8 }{11}\right)^n}\)

OpenStudy (solomonzelman):

what would this sum be?

OpenStudy (anonymous):

8/3 ?

OpenStudy (solomonzelman):

yes, very nice

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ \left(\frac{ 2 }{11}\right)^n+\sum_{ n=1 }^{ \infty } ~ \left(\frac{ 8 }{11}\right)^n}=\frac{2}{9}+\frac{8}{3}\)

OpenStudy (solomonzelman):

questions?

OpenStudy (anonymous):

no, thank you so much!! you have no idea how much clearer this is! Would you mind helping me with one last question? its different and a bit scary looking

OpenStudy (solomonzelman):

Yes, sure...

OpenStudy (anonymous):

thank you.. ill try to type it with the equation function option so its understandable

OpenStudy (solomonzelman):

or you can draw it if you want to:)

OpenStudy (anonymous):